# Uplift modeling with continuous treatments: A predict-then-optimize approach

### v1

@@ 5.2 Evaluation metrics
In the optimization step, several metrics evaluate how well the treatments are allocated, using the estimated CADEs, under the imposed constraints. The primary metric is Value ($V_{B}$), where a distinction is made between expected (Eq. 9), prescribed (Eq. 10), and optimal (Eq. 11) value. The cost-insensitive version where costs and benefits are equal over all entities is noted as $U_{B}$. Regret is defined as the difference between the full-information optimal value and the value achieved by the prescribed decision (Mandi et al., 2023). In other words, it measures how suboptimal the prescribed decision is compared to the optimal solution under ground-truth parameters. When CADRs are perfectly estimated, regret is zero. However, even with imperfect CADR estimations, regret can still be zero if the assignment policy remains optimal despite inaccuracies in the CADR estimates.
$Regret_{B}=V_{B}^{opt}-V_{B}^{presc}=V\left(\pi_{B}(\tau),\tau\right)-V\left(\pi_{B}(\hat{\tau}),\tau\right)$ (13)
@@
$RegretN_{B}=\frac{V_{B}^{opt}-V_{B}^{presc}}{V_{B}^{opt}}=\frac{V\left(\pi_{B}(\tau),\tau\right)-V\left(\pi_{B}(\hat{\tau}),\tau\right)}{V\left(\pi_{B}(\tau),\tau\right)}$ (14)
@@ 5.3 Results and discussion
In this section, we describe the experimental evaluation of our proposed predict-then-optimize framework, focusing on its ability to handle continuous treatments, optimize dose allocation under various constraints, and balance possibly conflicting objectives such as fairness and policy value. In all experiments, the number of bins $\delta$ is fixed to 10, motivated by C, and the level of confounding bias is constant. All experiments are implemented using Python 3.9 and can be reproduced with the code available on Github. All ILPs are solved using Gurobi 11.0.
@@ 5.3.1 Experiment 1: Performance of dose-response estimators
The first experiment aims to compare the performance of different dose-response estimators in terms of both prediction accuracy and capacity for good dose allocation. We evaluate four dose-response estimators: S-Learner (rf) and S-Learner (mlp) (Künzel et al., 2019), DRNet (Schwab et al., 2020), and VCNet (Nie et al., 2021). For each estimator, we conduct an internal 5-fold cross-validation loop, tuning hyperparameters across a grid search (details in D). The model with the lowest average MSE on factual outcomes over all folds is selected. In this experiment, we do not consider cost-sensitivity or fairness constraints. The budget is incrementally raised to observe the corresponding variation in $U^{presc}_{B}$. The results are displayed in Figure 3 and Table 3. The analysis provides two insights.
Figure 3: This figure shows $U^{presc}$ for four dose-response estimators as the available budget levels increase.
@@
S-Learner (rf) 0.059 0.869 0.941
S-Learner (mlp) 0.045 0.840 0.915
DRNet 0.037 0.909 0.935
VCNet 0.075 0.780 0.856
@@
First, in comparison to Uplift curves in a binary treatment setting, based on a ranking heuristic, the ILP with continuous treatments as input does not fully utilize the budget. Around a budget of 140, the ILP finds a cut-off point where it is more valuable to avoid overtreating certain entities. Unlike the heuristic approach which ranks treatment effects from high to low and selects the top $k$ within the budget, continuous treatment effects have a larger search space, allowing for partial treatments. This is evident in Figure 3 where the Uplift curves flatten around a budget of 140. Therefore, in Table 3, we include AUUC not only for the full budget (i.e., @250) but also for a partial budget of 140. Around this point, for all methods considered (and the full-information optimal solution), the policy value remains constant despite increasing budget.
Second, Table 3 shows a misalignment between the quality of CADR estimation (measured by MISE during prediction) and the quality of treatment allocation (measured by AUUC during the optimization step). MISE does not fully capture downstream task performance, as errors in critical areas of the curve have a greater impact on treatment allocation than those in less important regions. For instance, DRNet achieves the best MISE and AUUC@140 but ranks second in AUUC@250, behind the S-Learner (rf). Interestingly, the S-Learner (rf), which ranks third in MISE, achieves the highest AUUC@250. This suggests that the best inference methods do not always result in optimal decision-making, a notion similarly observed in cost-sensitive learning literature, where maximizing predictive accuracy does not necessarily yield the highest profit (Verbeke et al., 2012). Figure 9(a) illustrates this phenomenon: the S-Learner (rf) produces low-quality estimates for $S\in[0.4,0.6]$, resulting in a relatively high MISE. However, errors in this region have minimal impact on optimization since they correspond to low CADEs and are not selected anyway. Conversely, perfect treatment allocation does not require flawless predictions; effective allocation is possible even with imperfect dose-response estimates.
@@ 5.3.2 Experiment 2: Fairness trade-offs in treatment allocation
Figure 4: The effect of slack parameters $\epsilon_{DT}$ and $\epsilon_{DO}$ on respectively disparate treatment (allocation fairness) and disparate outcome (outcome fairness). Both panels use a logarithmic scale.
Figure 5: The x-axis represents the fairness constraint parameter for disparate treatment ($\epsilon_{DT}$), and the y-axis represents the fairness constraint parameter for disparate outcome ($\epsilon_{DO}$). Lower values of $\epsilon$ indicate stricter fairness constraints. $U^{presc}$ is normalized between 0 and 1 and is aggregated over multiple budgets (from 25 to 250 in increments of 25). The averaged normalized $U^{presc}$ is color-coded where blue means lower and yellow means higher
@@
The second experiment examines the trade-offs between policy value and business requirements as constraints, with a particular focus on fairness. Specifically, we examine both allocation fairness and outcome fairness, analyzing the impact of tightening or loosening these fairness constraints on overall policy value. In this experiment, we fix the budget and disregard cost-sensitivity. We use the S-Learner (rf) which performed best in terms of AUUC as the dose-response estimator. We vary two key fairness constraints: allocation fairness, which measures the disparity in assigned doses between groups, and outcome fairness, which reflects the difference in treatment outcomes. These constraints are regulated by two slack parameters: $\epsilon_{DT}$ for allocation fairness and $\epsilon_{DO}$ for outcome fairness.
@@
Figure 4(b) focuses on outcome fairness, influenced by the parameter $\epsilon_{DO}$. Unlike treatment fairness, discrepancies between estimates and ground truth emerge here. During optimization, outcome fairness is based on estimated CADEs, not the ground truth. While outcome fairness can be restricted by adjusting $\epsilon_{DO}$ (shown by the red line), this does not ensure a tight match with ground truth fairness (blue line). This gap depends on the quality of CADR estimates, highlighting the challenges of predictive analytics where fairness in predictions does not always translate into real-world fairness.
@@
The setup of Figure 6(c) is similar to the one in Figure 5 but also examines varying data-generating processes where the two groups defined by the protected feature are more different. This is controlled by the parameter $\gamma$, with higher values indicating greater differences in ground-truth average treatment effects (ATEs) between the two protected groups. Figure 6(c) plots the results for three increasing values of $\gamma$ (the case for $\gamma=0$ is displayed in Figure 5). As the difference in ATEs between the groups increases, the effect of slack parameters $\epsilon_{DT}$ and $\epsilon_{DO}$ on the objective value (normalized $U^{presc}$, averaged over different budgets) becomes more pronounced. Strict slack parameters are tolerable without significant value loss when the groups are similar ($\gamma=0$, Figure 5). In contrast, even mild slack parameters significantly affect the objective value when ATEs are most different ($\gamma=10$, Panel 6(c)).
@@ 5.3.3 Experiment 3: The impact of cost-sensitivity on utility
This experiment explores the effect of incorporating cost-sensitivity into policy optimization. Building upon the previous experiments, we now account for cost-sensitivity in the adapted objective function of ILP, considering instance-dependent treatment costs, $\mathbf{C}$, and outcome benefits, $\mathbf{b}$.
(a) Cost-insensitive $U_{B}^{presc}$ across increasing budget.
(b) Cost-sensitive $V_{B}^{presc}$ across increasing budget.
@@
This experiment highlights the importance of cost-sensitivity in optimizing treatment allocation under constrained budgets. By taking into account treatment costs and outcome benefits, cost-sensitive policies achieve higher values in cost-sensitive metrics, especially when resources are limited and not all entities can receive full treatment. These policies prioritize entities with the highest benefit-to-cost ratios, improving allocation efficiency. However, they underperform on cost-insensitive utility metrics, underlining the need for alignment between optimization objectives and specific decision-making goals. These findings are relevant for applications such as healthcare, lending, and human resource management, where treatment costs and benefits vary across entities, regardless of whether the goal is to maximize profitability or allocate scarce resources more efficiently. Furthermore, the modularity of the predict-then-optimize framework facilitates easy adaptation to changing objectives or constraints without requiring the retraining of predictive models. This makes it well-suited for dynamic environments where goals and resource availability evolve over time. An extension of this experiment could explore the inclusion of application-dependent and entity-dependent treatment cost functions and their interaction with the previously researched fairness constraints, adding complexity to the optimization process.
@@ 6 Conclusion, limitations, and further research
UM is extensively discussed in the literature, with applications such as lending, healthcare, HR, marketing, and maintenance. However, typically, no clear distinction between UM and CATE estimation is made, where CATE estimation should be viewed as a component of the broader UM field and is used as input for an optimization task. By clearly distinguishing these steps, our approach accommodates more complex treatments, as demonstrated in our focus on continuous treatments, while allowing flexible integration of constraints like fairness considerations, and cost-sensitive objectives. Our main contributions are (i) defining UM, (ii) extending it to handle continuous treatments with customizable constraints and objectives, (iii) defining fairness considerations as explicit ILP constraints, and (iv) demonstrating its capabilities through experiments. This framework’s flexibility enables the use of pre-trained models and avoids challenges in decision-focused learning by allowing constraint integration without retraining, particularly advantageous when, for example, unexpected operational constraints arise. Continuous treatments allow fine-grained dose-based interventions, better suited for various applications, as shown in Table 1.
@@
While effective, our framework has limitations. We do not focus on scalability or heuristic solutions, which are essential for large-scale applications, especially if the number of treatment bins would increase, but fall outside our current scope. Additionally, the assumption of deterministic costs simplifies the problem but does not account for stochastic or uncertain treatment costs, which are present in other settings (Haupt and Lessmann, 2022). Another limitation is the focus on single-phase treatments, where many real-world scenarios involve sequential decisions. We also address fairness at the group level, leaving individual-level fairness for future exploration. Finally, experiments are based on a single dataset and data-generating process, which may limit generalizability.
This work is the first step in defining and providing an initial solution for UM with continuous treatments. Future research could extend this work by exploring varied allocation schemes, treatment characteristics, or sequential decision-making using dynamic policy or reinforcement learning. In cases of time dependencies, treatments are administered multiple times over a period, or treatment effects change over time, dynamic policy learning or reinforcement learning methods could be employed to handle sequential treatments and long-term effects (Berrevoets et al., 2022). Extending the framework to accommodate multiple treatment types, whether mutually exclusive or not, would create new opportunities in situations where multiple intervention types are available (Zhan et al., 2024). Further research could investigate treatment effects in a network setting where spillover effects take place, making an entity’s outcomes dependent on others’ assigned doses and violating the stable unit treatment value assumption (Zhao et al., 2024). Additionally, while this study focuses on one-dimensional dose-based treatments, other applications may involve a high-dimensional treatment space (Berrevoets et al., 2020). When optimizing under uncertainty, incorporating conformal predictions, where CATE distributions or intervals around CADRs are estimated instead of point estimates, could better inform personalized decision-making (Vanderschueren et al., 2023a; Schröder et al., 2024). Finally, integrating DFL approaches, which, unlike our predict-then-optimize method, combine prediction and optimization in a single step, could potentially further improve policy outcomes by training models to directly make predictions that lead to better decisions (Mandi et al., 2023).

### v2

@@ 5.2 Evaluation metrics
In the optimization step, several metrics evaluate how well the treatments are allocated, using the estimated CADEs, under the imposed constraints. The primary metric is the total Value ($V$), computed as the sum of individual $V_{i}$ across all $N$ entities. We distinguish $V^{exp}$ (Eq. 19), $V^{presc}$ (Eq. 20), and $V^{opt}$ (Eq. 21) corresponding to the expected, prescribed, and optimal values, respectively. The cost-insensitive version, where costs and benefits are equal over all entities, is noted as $\stackrel{{\scriptstyle}}{{U}}$. Regret is defined as the difference between the full-information optimal value and the value achieved by the prescribed decision (Mandi et al., 2023). In other words, it measures how suboptimal the prescribed decision is compared to the optimal solution under ground-truth parameters. When CADRs are perfectly estimated, regret is zero. However, even with imperfect CADR estimations, regret can still be zero if the assignment policy remains optimal despite inaccuracies in the CADR estimates.
$Regret=V^{opt}-V^{presc}=V\left(\stackrel{{\scriptstyle B}}{{\Pi}}(\tau),\tau\right)-V\left(\Pi(\hat{\tau}),\tau\right)$ (23)
@@
$RegretN_{B}=\frac{\stackrel{{\scriptstyle}}{{V}}^{opt}-\stackrel{{\scriptstyle}}{{V}}^{presc}}{\stackrel{{\scriptstyle}}{{V}}^{opt}}=\frac{V\left(\stackrel{{\scriptstyle B}}{{\Pi}}(\tau),\tau\right)-V\left(\stackrel{{\scriptstyle B}}{{\Pi}}(\hat{\tau}),\tau\right)}{V\left(\stackrel{{\scriptstyle B}}{{\Pi}}(\tau),\tau\right)}$ (24)
@@ 5.3 Results and discussion
In this section, we describe the experimental evaluation of our proposed predict-then-optimize framework, focusing on its ability to handle continuous treatments, optimize dose allocation under various constraints, and balance possibly conflicting objectives such as fairness and policy value. In all experiments, the number of bins $\delta$ is fixed to 10, motivated by D, and the level of confounding bias is constant. All experiments are implemented using Python 3.9 and can be reproduced with the code available on Github. All ILPs are solved using Gurobi 11.0. In all experimental scenarios presented, the ILP solver converged, guaranteeing optimal solutions for the given problem instances. Solver convergence was systematically verified by checking the solver status after each optimization run.
@@ 5.3.1 Experiment 1: Performance of dose-response estimators
The first experiment aims to compare the performance of different dose-response estimators in terms of both prediction accuracy and capacity for good dose allocation. We evaluate four dose-response estimators: S-Learner (rf) and S-Learner (mlp) (Künzel et al., 2019), DRNet (Schwab et al., 2020), and VCNet (Nie et al., 2021). For each estimator, we conduct an internal 5-fold cross-validation loop, tuning hyperparameters across a grid search (details in E). The model with the lowest average MSE on factual outcomes over all folds is selected. In this experiment, we do not consider cost-sensitivity or fairness constraints. The budget is incrementally raised to observe the corresponding variation in $U^{presc}_{B}$. Each model’s estimated CADRs are visualized in F.1.
We allocate treatments using two distinct methods. The first approach utilizes the ILP introduced in Section 4.2. The second employs a heuristic inspired by the multi-treatment framework (Olaya et al., 2020). Here, each discrete dose is treated as a separate intervention and, formally, the optimal treatment for an entity with features $\mathbf{x}_{i}$ is $\operatorname{argmax}(\mathbf{\hat{\tau}}(\mathbf{x}_{i}))$. Entities are then ranked according to the uplift of their optimal treatment, and doses are assigned greedily in that order until the budget is depleted. The results are displayed in Figure 3 and Table 3. The analysis provides two insights.
Figure 3: This figure shows $U^{presc}$ for four dose-response estimators as the available budget levels increase. Doses are allocated with an ILP and a heuristic approach.
@@
Heur. ILP Heur. ILP
S-Learner (rf) 0.060 0.837 0.892 0.926 0.951
S-Learner (mlp) 0.044 0.813 0.829 0.877 0.884
DRNet 0.049 0.649 0.799 0.729 0.797
VCNet 0.055 0.827 0.827 0.922 0.922
@@
First, compared to typical uplift curves from binary treatment scenarios (based on a ranking heuristic), both the multi-treatment heuristic and the ILP with continuous treatments as input do not fully utilize the budget. Around a budget of 140, they find a point where further treatment becomes counterproductive —i.e., it is more valuable to avoid overtreating certain entities— due to the non-monotonic nature of CADRs. From a budget of 140 onwards, per estimator, the heuristic and ILP approach find the same solution (i.e., selection of the dose with maximum estimated CADE) and they converge in terms of $U^{presc}$. Unlike the heuristic approach, which ranks binary treatment effects from high to low and selects the top $k$ within the budget, continuous treatment effects have a larger search space, allowing for partial treatments. This is evident in Figure 3 where the Uplift curves flatten around a budget of 140. Therefore, in Table 3, we include AUUC not only for the full budget (i.e., @250) but also for a partial budget of 140. Around this point, for all methods considered (and the full-information optimal solution), the policy value remains constant despite increasing budget.
Second, Table 3 shows a misalignment between the quality of CADR estimation (measured by MISE during prediction) and the quality of treatment allocation (measured by AUUC during the optimization step). MISE does not fully capture downstream task performance, as errors in critical areas of the curve have a greater impact on treatment allocation than those in less important regions — which holds true for both heuristic and the ILP approaches.
For example, although S-Learner (mlp) has the lowest MISE, its final allocations rank second or third. DRNet shows the second-best MISE but performs worst in terms of AUUC at both budget points, irrespective of the treatment assignment method. Conversely, S-Learner (rf), despite its high MISE, outperforms all others in uplift effectiveness. This suggests that the best inference methods do not always result in optimal decision-making, a notion similarly observed in cost-sensitive learning literature, where maximizing predictive accuracy does not necessarily yield the highest profit (Verbeke et al., 2012). Figure 9(a) illustrates this phenomenon: the S-Learner (rf) produces low-quality estimates for $S\in[0.4,0.6]$, resulting in a relatively high MISE. However, errors in this region have minimal impact on optimization since they correspond to low CADEs and are not selected anyway. DRNet, on the other hand, has relatively small errors overall but struggles significantly in the crucial high-impact region of $S\in[0.9,1.0]$, where the ground-truth CADR sharply declines (see Figure 9(c)). Thus, accurate dose allocation doesn’t strictly require perfect predictions, provided the predictions capture the critical areas effectively. Conversely, perfect treatment allocation does not require flawless predictions; effective allocation is possible even with imperfect dose-response estimates. A further analysis of runtime for increasing problem sizes is provided in F.2. While the heuristic approach scales well for increasing problem sizes, it does not allow additional side constraints like fairness, which are allowed for the ILP.
@@ 5.3.2 Experiment 2: Fairness trade-offs in treatment allocation
Figure 4: The effect of slack parameters $\epsilon_{DT}$ and $\epsilon_{DO}$ on, respectively, disparate treatment (allocation fairness) and disparate outcome (outcome fairness). Both panels use a logarithmic scale.
Figure 5: The x-axis represents the fairness constraint parameter for disparate treatment ($\epsilon_{DT}$), and the y-axis represents the fairness constraint parameter for disparate outcome ($\epsilon_{DO}$). Lower values of $\epsilon$ indicate stricter fairness constraints. $U^{presc}$ is normalized between 0 and 1 and is aggregated over multiple budgets (from 25 to 250 in increments of 25). The averaged normalized $U^{presc}$ is color-coded, where blue means lower and yellow means higher.
@@
The second experiment examines the trade-offs between policy value and business requirements as constraints, with a particular focus on fairness. Specifically, we examine both allocation fairness and outcome fairness, analyzing the impact of tightening or loosening these fairness constraints on overall policy value. In this experiment, we fix the budget and disregard cost-sensitivity. We use the S-Learner (rf), which performed best in terms of AUUC, as the dose-response estimator. We vary two key fairness constraints: allocation fairness, which measures the disparity in assigned doses between groups, and outcome fairness, which reflects the difference in treatment outcomes. These constraints are regulated by two slack parameters: $\epsilon_{DT}$ for allocation fairness and $\epsilon_{DO}$ for outcome fairness.
@@
Figure 4(b) examines outcome fairness as regulated by the parameter $\epsilon_{DO}$, revealing a distinct challenge compared to treatment fairness: a discrepancy between estimated and ground truth fairness. This divergence arises because optimization is based on estimated CADEs rather than true values. Although tightening $\epsilon_{DO}$ (illustrated by the red line) can restrict estimated disparities, it does not guarantee alignment with ground truth fairness (blue line). This misalignment is driven by the quality of the CADE estimates and underscores a broader challenge in predictive analytics—namely, that fairness in predictions does not always translate into fairness in actual outcomes. Nevertheless, we argue that incorporating constraints on expected outcomes across groups remains a principled and ethically sound design choice. Even when individual-level outcomes are uncertain, they still provide the most informative basis available for assessing how different groups may benefit from the allocation policy.
@@
The setup of Figure 6(c) is similar to the one in Figure 5, but also examines varying data-generating processes where the two groups defined by the protected feature are more different. This is controlled by the parameter $\gamma$, with higher values indicating greater differences in ground-truth average treatment effects (ATEs) between the two protected groups. Figure 6(c) plots the results for three increasing values of $\gamma$ (the case for $\gamma=0$ is displayed in Figure 5). As the difference in ATEs between the groups increases, the effect of slack parameters $\epsilon_{DT}$ and $\epsilon_{DO}$ on the objective value (normalized $U^{presc}$, averaged over different budgets) becomes more pronounced. Strict slack parameters are tolerable without significant value loss when the groups are similar ($\gamma=0$, Figure 5). In contrast, even mild slack parameters significantly affect the objective value when ATEs are most different ($\gamma=10$, Panel 6(c)).
@@ 5.3.3 Experiment 3: The impact of cost-sensitivity on utility
In many applications, the ultimate goal is profit or cost reduction, not just uplift. This experiment demonstrates that optimizing for the true objective (value) yields different policies than optimizing for uplift alone. This experiment explores the effect of incorporating cost-sensitivity into policy optimization. Building upon the previous experiments, we now account for cost-sensitivity in the adapted objective function of ILP, considering instance-dependent treatment costs, $\mathbf{C}$, and outcome benefits, $\mathbf{b}$.
(a) Cost-insensitive $\stackrel{{\scriptstyle}}{{U}}^{presc}$ across increasing budget.
(b) Cost-sensitive $\stackrel{{\scriptstyle}}{{V}}^{presc}$ across increasing budget.
@@
This experiment highlights the importance of cost-sensitivity in optimizing treatment allocation under constrained budgets. By taking into account treatment costs and outcome benefits, cost-sensitive policies achieve higher values in cost-sensitive metrics, especially when resources are limited and not all entities can receive full treatment. These policies prioritize entities with the highest benefit-to-cost ratios, improving allocation efficiency. However, they underperform on cost-insensitive utility metrics, underlining the need for alignment between optimization objectives and specific decision-making goals. The cost-sensitive policy, as expected, sacrifices some uplift if it leads to higher net value. We see differences in who gets treated: for example, some individuals with high uplift but low benefit may not be treated under the value-based policy, whereas they would under an uplift-only policy These findings are relevant for applications such as healthcare, lending, and human resource management, where treatment costs and benefits vary across entities, regardless of whether the goal is to maximize profitability or allocate scarce resources more efficiently. Furthermore, the modularity of the predict-then-optimize framework facilitates easy adaptation to changing objectives or constraints without requiring the retraining of predictive models. This makes it well-suited for dynamic environments where goals and resource availability evolve over time. An extension of this experiment could explore the inclusion of application-dependent and entity-dependent treatment cost functions and their interaction with the previously researched fairness constraints, adding complexity to the optimization process.
@@ 6 Conclusion, limitations, and further research
UM is extensively discussed in the literature, with applications such as lending, healthcare, HR, marketing, and maintenance. However, typically, no clear distinction between UM and CATE estimation is made, where CATE estimation should be viewed as a component of the broader UM field and is used as input for an optimization task. By clearly distinguishing these steps, our approach accommodates more complex treatments, as demonstrated in our focus on continuous treatments, while allowing flexible integration of constraints like fairness considerations and cost-sensitive objectives. Our main contributions are (i) defining UM, (ii) extending it to handle continuous treatments with customizable constraints and objectives, (iii) defining fairness considerations as explicit ILP constraints, and (iv) demonstrating its capabilities through experiments. This framework’s flexibility enables the use of pre-trained models and avoids challenges in decision-focused learning by allowing constraint integration without retraining, particularly advantageous when, for example, unexpected operational constraints arise. Continuous treatments allow fine-grained dose-based interventions, better suited for various applications, as shown in Table 1.
@@
While effective, our framework has limitations. The ILP approach suffers from limited scalability. In contrast, the proposed heuristic handles larger problem sizes more efficiently. However, it may yield suboptimal solutions and does not account for side constraints such as fairness. Additionally, the assumption of deterministic costs simplifies the problem but does not account for stochastic or uncertain treatment costs, which are present in other settings (Haupt and Lessmann, 2022). Another limitation is the focus on single-phase treatments, where many real-world scenarios involve sequential decisions. We also address fairness at the group level, leaving individual-level fairness for future exploration. Finally, experiments are based on a single dataset and data-generating process, which may limit generalizability.
This work is the first step in defining and providing an initial solution for UM with continuous treatments. Future research could extend this work by exploring varied allocation schemes, treatment characteristics, or sequential decision-making using dynamic policy or reinforcement learning. In cases of time dependencies, treatments are administered multiple times over a period, or treatment effects change over time, dynamic policy learning or reinforcement learning methods could be employed to handle sequential treatments and long-term effects (Berrevoets et al., 2022). Extending the framework to accommodate multiple treatment types, whether mutually exclusive or not, would create new opportunities in situations where multiple intervention types are available (Zhan et al., 2024). Further research could investigate treatment effects in a network setting where spillover effects take place, making an entity’s outcomes dependent on others’ assigned doses and violating the stable unit treatment value assumption (Zhao et al., 2024). Additionally, while this study focuses on one-dimensional dose-based treatments, other applications may involve a high-dimensional treatment space (Berrevoets et al., 2020). When optimizing under uncertainty, incorporating conformal predictions, where CATE distributions or intervals around CADRs are estimated instead of point estimates, could better inform personalized decision-making (Vanderschueren et al., 2023a; Schröder et al., 2024). Finally, integrating DFL approaches, which, unlike our predict-then-optimize method, combine prediction and optimization in a single step, could potentially further improve policy outcomes by training models to directly make predictions that lead to better decisions (Mandi et al., 2023). One could also explore hybrid approaches –– for example, incorporating fairness constraints directly into model training (a partial DFL approach) while still performing an optimization step for other parameters or constraints. Such methods might alleviate some of the burden on the post-prediction optimization